OpenStudy (kayders1997):

Please help I'm stuck but I have a start find the integral of sinx/1+cos^2x would anyone turn the numerator into sin^2x?

OpenStudy (kayders1997):

@zepdrix @ganeshie8 @vishweshshrimali5

OpenStudy (kayders1997):

Denomiator*

OpenStudy (anonymous):

try \[u=\cos(x)\]should work immediately

OpenStudy (kayders1997):

I did u=1+cos^2x probably why I'm having trouble

OpenStudy (anonymous):

\[\int \frac{\sin(x)dx}{1+\cos^2(x)}\]right ?

OpenStudy (kayders1997):

Yes

OpenStudy (anonymous):

you get \[-\int \frac{1}{1+u^2}du\]right a way , which should look like a familiar derivative

OpenStudy (anonymous):

think of a well know function whose derivative is \[\frac{1}{1+x^2}\]

OpenStudy (kayders1997):

Inverse tangent? I think

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

believe it or not, with the minus sign included it is actually the derivative of arccotangent, but no one ever used that

OpenStudy (anonymous):

*uses

OpenStudy (kayders1997):

Yeah, interesting so how would I know when to just use cos instead of cos^2x+1?

OpenStudy (anonymous):

practice

OpenStudy (anonymous):

plus if you put \[u=1+\cos^2(x)\] then you get \[du=-2\cos(x)\sin(x)\] which is no where in sight

OpenStudy (kayders1997):

I was going to say when I did it that's probably why I got stuck because I didn't know whAt to do after I found the du= part

OpenStudy (kayders1997):

Right

OpenStudy (anonymous):

it is like learning how to factor just as tedious, just as boring, and for the most part just as useless, although some people dig it

OpenStudy (kayders1997):

Lol yes :D

OpenStudy (anonymous):

i would not include myself in the group that finds it at all interesting

OpenStudy (kayders1997):

It's okay, thank you for your help :)

OpenStudy (anonymous):

yw

OpenStudy (kayders1997):

:)